OpenStudy (anonymous):

Angle S and angle T are congruent. Angle S measures 2x + 5 and angle T measures 5x – 25. What is the measure of angle S?

OpenStudy (bahrom7893):

What do congruent angles look like?

OpenStudy (anonymous):

theres no picture thats all it says

OpenStudy (bahrom7893):

Well, look up congruent angles on google. You'll get a ton of pics.

OpenStudy (anonymous):

angle A = 2x + 5 angle T = 5x - 25 2x + 5 = 5x - 25 2x - 5x = -25 - 5 -3x = -30 x = 10 Angle A = 2x + 5 : 2(10) + 5 = 20 + 5 = 25 Angle B = 5x - 25 : 5(10) - 25 = 50 - 25 = 25 congruent angles are equal, and both these angles are 25

OpenStudy (bahrom7893):

Kellie, please don't give away the answer, as you could tell, I was trying to get him/her to learn this -_-.

OpenStudy (anonymous):

I am sorry...I didn't even know you were on...I was busy typing....again,I am sorry